OpenStudy (anonymous):

solve [\lim_{x \rightarrow 1} [(1+3x)/(1+4x^2+3x^4)]^3\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1}[(1+3x)/(1+4x^2+3x^4)]^3\]

OpenStudy (anonymous):

Just plug in 1 for x..

OpenStudy (anonymous):

yes and i got {(1+3)/(1+4+3)]^3 now how do i do the third power

OpenStudy (anonymous):

\[1+3 = 4\]\[1 + 4 + 3 = 8\]\[{4\over 8} = {1 \over 2}\] \[\implies [{1+3 \over 1+4 + 3}]^3 = [{4\over 8}]^3 = [{1\over 2}]^3 = {1^3 \over 2^3} = {1\over 8}\]